Using the letters A,M,O,S,U. We can form 120 five letter "words". If these "words" are arranged in alphabetical order, then what position the "word"
USAMO occupies?

Answer by josmiceli(6793)

I come up with the 115th position in the list.
The list of 120 words is made up of:
24 words that start with "A"
24 words that start with "M"
24 words that start with "O"
24 words that start with "S"
24 words that start with "U"
5 X 24 = 120
"USAMO" occurs after 4 X 24 words, or after 96 words on the list
the 97th word is "UAMOS"
6 X 4 = 24 words that start with "U"
There are 6 words that start with "US"
There are 6 words that start with "UA"
There are 6 words that start with "UM"
There are 6 words that start with "UO"
The first word that starts with "US" is "USAMO"
So, the number on the list is:
96 + (3*6) + 1 = 96 + 19 = 115